Please help me solve this:
Mike's house payments rose from $295 to $310 per month. In the past 12 months, he paid $3615. How many payments at each rate did he make?

Answer by ankor@dixie-net.com(22740)

Mike's house payments rose from $295 to $310 per month. In the past 12 months, he paid $3615. How many payments at each rate did he make?
:
Let x = number of $310 payments
Let (12-x) = number of $295 payments
:
$310 payments + $295 payments = 3615
310x + 295(12-x) = 3615
310x + 3540 - 295x = 3615
310x - 295x = 3615 - 3540
15x = 75
x = 75/15
x = 5 ea $310 payments
:
12 - 5 = 7 ea $295 payments
:
:
Check: 5(310) + 7(295) = 3615